# DB-GPT: Empowering Database Interactions with Private Large Language Models

### 29 Dec 2023 v1

@@ 2.2 Deploy and Inference: Service-oriented Multi-model Framework
SMMF consists of two principal components, namely the model inference layer and the model deployment layer. Specifically, the model inference layer is designed for accommodating various LLM inference platforms, including vLLM (Kwon et al., 2023), Text Generation Inference (TGI) (Huggingface, 2021), and TensorRT (NVIDIA, 2021). The model deployment layer serves as an intermediary between the underlying inference layer and the upper-level model serving functionalities.
@@ 3.2 Encoder in RAG
For demo and testing purposes, our system is deployed on a server on Alibaba Cloud with 30G RAM, a 8 logical cores (Intel Xeon(Ice Lake) Platinum 8369B), and a NVIDIA A100 80G Tensor Core GPU.
@@ 4.3 SMMF Evaluation
Clarified in Section 2.2, our DB-GPT integrates FastChat as the main inference framework.
@@
The test is performed on a server with 90G RAM, a 12 logical cores (Intel Xeon(Ice Lake) Platinum 8352V) CPU, and a RTX 4090 Tensor Core GPU with 24G GPU RAM. Across all the experiments, we use the same prompt with $8$ tokens as the input while setting the output length to be $256$ tokens.
@@
Seen in Table 5 and Table 6, the results indicate that the use of the FastChat framework for model inference significantly increases the throughput of the model, while substantially reducing both the first token latency and the overall inference latency. Moreover, as the number of concurrent users rises, the performance improvements gained from utilizing the FastChat framework for inference become particularly pronounced. As a result, DB-GPT choose to integrate FastChat as the default inference framework used for SMMF.
@@
FTL
IL
Throughput
Qwen-7b
@@
FTL
IL
Throughput
Baichuan-7b

### 03 Jan 2024 v2

@@ 2.2 Deploy and Inference: Service-oriented Multi-model Framework
SMMF consists of two principal components, namely the model inference layer and the model deployment layer. Specifically, the model inference layer is designed for accommodating various LLM inference platforms, including vLLM (Kwon et al., 2023), HuggingFace Transformers (HF) (hf), Text Generation Inference (TGI) (Huggingface, 2021), and TensorRT (NVIDIA, 2021). The model deployment layer serves as an intermediary between the underlying inference layer and the upper-level model serving functionalities.
@@ 3.2 Encoder in RAG
For demo and testing purposes, unless otherwise specified, our system is deployed on a server on Alibaba Cloud with 30G RAM, a 8 logical cores (Intel Xeon(Ice Lake) Platinum 8369B), and a NVIDIA A100 80G Tensor Core GPU.
@@ 4.3 SMMF Evaluation
Clarified in Section 2.2, our DB-GPT integrates vLLM as the main inference framework.
@@
The test is performed on a server with 629G RAM, a hard drive with 1TB HDD, a 40 logical cores (Intel Xeon Processor (Skylake, IBRS)) CPU at 2992.953MHz, and a NVIDIA A100-PCIE GPU with 40G GPU RAM. Across all the experiments, we use the same prompt with $8$ tokens as the input while setting the output length to be $256$ tokens.
@@
Seen in Table 5 and Table 6, the results indicate that the use of the vLLM framework for model inference significantly increases the throughput of the model, while substantially reducing both the first token latency and the overall inference latency. Moreover, as the number of concurrent users rises, the performance improvements gained from utilizing the vLLM framework for inference become particularly pronounced. As a result, DB-GPT choose to integrate vLLM as the default inference framework used for SMMF.
@@
FTL(ms)
IL(s)
Throughput (tokens)
Qwen-7b-Chat
@@
FTL (ms)
IL (s)
Throughput (tokens)
Baichuan-7b-Chat